# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ 4.3.2 Klein-Nishina scattering in high energies
Using the relation $F_{\gamma}/n_{\gamma}=T_{\gamma}/3.7$ and the KN cross section (4.1) that depends on drifting electrons’ $\gamma_{e}$, the variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path is given by
@@
in the contrast with Eq. (4.7) for the Thomason scattering case.
@@
$\frac{dN_{e}}{dr}$ $=$ $-\frac{1}{\xi^{\rm KN}_{e}}N_{e}=-n_{\gamma}\sigma^{e}_{{}_{\rm KN}}N_{e},$ (4.13)
which recasts as
@@
Here we approximate $T_{\gamma}$ by the characteristic temperature $T^{+}_{\gamma}$ for the same argument discussed in Eqs. (4.5) and (4.10). Figure 5 gives the numerical integration (4.15) that shows the non-trivial number of drifting electrons at ultra-high energies, which is in distinct contrast with the completely negligible one in the Thomason scattering case (4.10). This non-trivial number $N_{e}$ of drifting electrons ($\gamma_{e}\gg 1$) means a superthermal spectrum in the high-energy tail.
Figure 5: We show the fraction $N_{e}/\bar{N}_{e}$ (4.15) of accelerated electrons drifting out of bulk electrons at different temperatures $T^{+}_{\gamma}\gg m_{e}$ near the horizon. The $N_{e}/\bar{N}_{e}$ value is exponentially sensitive to the horizon temperature $T_{\gamma}^{+}$. The probability of accelerated electrons reaching high energies is small for temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, and the Compton-rocket effect and runaway dynamics might not be observationally relevant. These figures also present the proton case by substituting the subscript $e\rightarrow p$.
Figure 5 shows that for Lorentz factors $\gamma_{e}<10^{3}$, (a) the accelerated electron fraction $N_{e}/\bar{N}_{e}$ decreases as $\gamma_{e}$ increases, which may approximate as a decay power law $(\gamma_{e})^{-\nu}$ ($\nu>0$); (b) the fraction approach a constancy $(\gamma_{e})^{0}$ for $\gamma_{e}>10^{3}$:
$(a)~\frac{N_{e}}{\bar{N}_{e}}\propto(\gamma_{e})^{-\nu},~(\gamma_{e}<10^{3});~~~(b)~\frac{N_{e}}{\bar{N}_{e}}~\propto(\gamma_{e})^{0},~(\gamma_{e}>10^{3}).$ (4.16)
The feature (a) is not unexpected in a usual physical situation where the probability of drifting electrons running to high energies becomes smaller. However, the feature (b) is peculiar, showing the high-energy superthermal spectrum tail becomes very flat when $\gamma_{e}>10^{3}$. It is the exact attributes of unstable runaway behaviour. Since the KN cross-section decreases as the drifting electrons’ energies increase, they continue accelerating to higher and higher energies by the Compton-rocket effect without dissipating much energy into photons and pairs. Such an avalanche dynamics is crucially important for producing UHE particles of Lorentz factors $\gamma_{e}\gg 1$.
Using $\Delta\gamma^{\rm KN}_{e}$ (4.12) and $(N_{e}/\bar{N}_{e})$ (4.15), we approximately calculate the asymptotic constancy for large $\gamma_{e}>10^{3}$
@@
which is exponentially suppressed by small horizon temperatures $T^{+}_{\gamma}/m_{e}$, as can be seen in Figs. 4 and 5. The excess of large $\gamma_{e}$ electrons $N_{e}$ in the superthermal tail deviating from the blackbody radiation spectrum in the case of the Sun ($T_{\odot}/m_{e}\ll 1$) is too small to be relevant for observations. However, the probability of $\gamma_{e}\gg 1$ electrons’ excess may be relevant for the detections by laser experiments for the physics of strong radiation fields in future.
For the case of electron-proton plasma interacting with high-energy photons $\epsilon_{\gamma}\sim m_{e}$, Equations of types (4.8,4.9) or (4.13,4.14) were used in the numerical studies Ref. [11] to discuss the superthermal spectrum tail of runaway electrons $N_{e}$ drifting backwards to massive protons due to the electric field induced by the charge separation between electrons and protons. Instead, we are in a situation where fully ionised electrons and protons are free, not in electron-proton plasma oscillations. They interact with photons and pairs at very high temperatures $T_{\gamma}\gg m_{e}$, energy density $\rho_{\gamma}\gg m_{e}^{4}$ and number density $n_{\gamma}\gg m_{e}^{3}$.
4.4 Runaway protons pulled by runaway electrons
A non-trivial amount of high-energy electrons running away indicates that we cannot ignore protons in the Compton-rocket effect on electrons, because of the electric field $E$ developed between separated electrons and protons. It is due to the very different cross sections $\sigma_{{}_{T}}\gg\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}$ of photons scattering with electrons and protons. The Compton-rocket effect on protons ($p_{p}=m_{p}\beta_{p}\gamma_{p}$)
@@ 4.4 Runaway protons pulled by runaway electrons
is much smaller than the one (3.2) on electrons.
@@
It is a highly non-linear dynamical system. We are interested in the situation when photons and pairs’ radiation force $\sigma_{{}_{T}}F_{\gamma}\propto T_{\gamma}^{4}(\beta_{e}\gamma_{e})^{2}$ in Eq. (4.20) is so strong to accelerate a significantly non-trivial fraction $N_{e}/\bar{N}_{e}$ of electrons drifting outward $dp_{e}/d\tau>0$ (4.20), and induce an electric field $E_{pe}=eN_{e}/(4\pi r^{2})$ to pull massive protons $dp_{p}/d\tau\approx E_{pe}>0$ (4.21) accelerating outwards in the radial direction, following the accelerated electrons. We assume that such a dynamical configuration can be realised by a large temperature $T_{\gamma}\gg m_{e}$, a significantly non-trivial fraction $N_{e}/\bar{N}_{e}$ and a large Lorentz factor $\gamma_{e}$. It is necessary to verify it using numerical simulations, under what conditions to realise such a dynamical configuration. Since protons are massive, if the photons and pairs temperature $T_{\gamma}$ is not large enough, instead of accelerating protons, the bulk electrons $\bar{N}_{e}$ are accelerated backwards by the induced electric field $E_{pe}$, as numerically studied in Ref. [11].
In this speculated configuration, we set that the bulk and drifting electron numbers are identical to the bulk and drifting proton numbers
@@
because of the global charge neutrality and conservation. The combination of Eqs. (4.20) and (4.21) yields the acceleration equation for electrons and protons together
@@
neglecting the radiation force on protons. Under the condition of a strong enough radiation flux $F_{\gamma}$ and the electric field $E$, we further model that drifting electrons $N_{e}$ and protons $N_{p}$ are moving together
@@
through the fluid of photons and pairs. The acceleration equation (4.23) is simplified as
$(m_{e}+m_{p})\frac{d(\beta_{e}\gamma_{e})}{d\tau}$ $\approx$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1],$ (4.25)
which is the same as the electron acceleration equation (3.2) or (4.3) by introducing an effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}\approx m_{p}$. Similarly, the combination of (3.1) and (4.18) yields the energy-gain equation for accelerated electrons and protons
@@
In the one-dimensional model, Equations (4.25) and (4.26) qualitatively describe an accelerating system of electrons and protons driven by the radiation force, assuming electrons and protons associated by their own electric field and accelerated to the approximately same Lorentz factor $\gamma_{e}\approx\gamma_{p}$. In this approximation, all dynamical equations remain the same, except for replacing the electron mass $m_{e}$ with the effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$ for a heavy system of electrons and protons. As previously anticipated, the effective mass $m^{\rm eff}_{e}\gg m_{e}$ replaces $m_{e}$ in the dynamic equations (4.3) and (4.25). As a consequence, the large temperature $T_{\gamma}$ radiation flux $F_{\gamma}$ acting on an electron is required to accelerate a heavy system of both electrons and protons.
@@
In addition, when electrons and protons drift through the fluid of photons and pairs, they undergo Klein-Nishina scattering processes. The KN cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons is much smaller than the one $\sigma^{e}_{{}_{\rm KN}}$ (4.1) of drifting electrons, i.e., $\sigma^{p}_{{}_{\rm KN}}\sim{\mathcal{O}}[(m_{p}/m_{e})^{2}]\sigma^{e}_{{}_{\rm KN}}$. Therefore, we neglect the cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons. In these approximations, all analyses and results in the previous sections remain the same. Both drifting electrons and protons run away, becoming UHE particles. Figures 5 also represent the fraction of accelerated protons with the substitution $e\rightarrow p$. In the following, without further specification, all discussions and calculations apply to a bound system of electrons and protons with an effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$. Note that for $\gamma_{e}\approx\gamma_{p}$ the accelerated protons’ energy $\gamma_{p}m_{p}$ is ${\mathcal{O}}[(m_{p}/m_{e})]$ larger than the accelerated electrons’ energy $\gamma_{e}m_{e}$.
These speculations and approximations in the one-dimensional model are used only to qualitatively gain some insights into the Compton-rocket accelerations of electrons and protons in an opaque fluid of photons and pairs. Numerical simulations are necessary.
5 Luminosities of UHE particles and blackbody radiation
@@
The accelerated electrons and protons reach ultra-high energies and stream out of the halo, carrying away the halo energy of photons and pairs. We calculate their luminosity and time variation, which can be relevant for observations. Assuming that the halo energy is large and the luminosity is small, we regard the halo as a steady state, like an energy reservoir. The thermal equilibrium is approximately maintained. The energy transfer and temperature cooling are adiabatic. The halo decays and shrinks slowly.
@@ 5.1 Blackbody radiation luminosity and MeV annihilation line
Due to the large opacity and temperature gradient in the centre of the fluid halo ($r_{+}<r<2r_{+}$), internal photons diffuse outwards in the radial direction. At the fluid halo outer boundary $r=2r_{+}$, where the temperature is about $2m_{e}$ and pair annihilation occurs, it becomes optically thin, and photons stream out as a continuous blackbody radiation of the characteristic temperature $2m_{e}$. One obtains the blackbody luminosity by solving the radiation transfer (photon diffusion) equation [31],
@@
where the photon luminosity $L^{+}_{\gamma}=\eta L_{\rm pl}$ (2.4) and mean-free path $\xi_{\gamma}\sim 10^{-4}\lambda_{e}$ (Fig. 3) are inside the centre of the fluid halo of photons and pairs. In this case, the small screen factor $\xi_{\gamma}/2r_{+}\sim 8.3\times 10^{-22}$ is small for the halo size $\sim 2r_{+}$ (2.3). The blackbody radiation luminosity (5.1) is a constant $L^{0}_{\gamma}\sim 10^{-22}L_{\rm pl}$, and the time scale (5.2) is large. These values depend on the halo parameters (2.3), (2.4) and (2.9).
This emission is analogous to the blackbody radiation from the Sun’s surface. To compare and contrast, we mention the Sun’s surface halo of the temperature $\sim 6\times 10^{3}C^{\circ}\approx 0.55$ eV and the radius $6.96\times 10^{10}$cm, where electrons are not relativistic ($\langle\gamma_{e}\rangle\approx 1$) and their densities are much smaller than the horizon halo under discussion. The Sun’s luminosity is $L_{\odot}=3.83\times 10^{33}{\rm ergs/sec}\approx 10^{-27}L_{\rm pl}$. Thermal photons diffuse from the Sun’s core at the temperature $T_{\odot}=15\times 10^{6}C^{0}=1.5\times 10^{4}{\rm eV}=3\times 10^{-2}m_{e}$. While in the horizon halo case, the blackbody has a characteristic temperature $\sim 2m_{e}$. Moreover, the radiation spectrum contains the electron-positron annihilation line and reheating of temperature increment $(11/4)^{1/3}$ before and after annihilation around $2m_{e}\approx 1$ MeV [21].
Figure 6: For the purpose of illustrating the spectral behaviour of UHE particle luminosity, we select the horizon temperature $T_{\gamma}^{+}=10m_{e},20m_{e}$ and acceleration number $A=10^{-3}$, to plot high-energy electron and proton luminosities $L_{e}/L_{\rm pl}$ (5.3) in terms of their Lorentz factor $\gamma_{e}$.

### v2

@@ 4.5 Unstable case: Klein-Nishina scattering at high energies
Using the relation $F_{\gamma}/n_{\gamma}=T_{\gamma}/3.7$ and the KN cross section (4.1) that depends on drifting electrons’ energy, i.e., Lorentz factor $\gamma_{e}$, we obtain the variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path given by
@@
in the contrast with $\Delta\gamma^{\rm T}_{e}$ (4.7) for the Thomason scattering case.
@@
$\frac{dN_{e}}{dr}$ $=$ $-\frac{1}{\xi^{\rm KN}_{e}}N_{e}=-n_{\gamma}\sigma^{e}_{{}_{\rm KN}}N_{e}.$ (4.13)
This rate equation is recast as
@@
Here, we approximate the radial-dependent temperature $T_{\gamma}(r)$ by the characteristic temperature $T^{+}_{\gamma}$ for the same arguments discussed in Eqs. (4.5) and (4.10). Figure 5 gives the numerical integration (4.15) that shows the non-trivial fraction $N_{e}/\bar{N}_{e}$ of drifting electrons at ultra-high energies, which is in distinct contrast with the completely negligible one (4.10) in the Thomson scattering case. This non-trivial number $N_{e}$ of drifting electrons ($\gamma_{e}\gg 1$) means a significant superthermal spectrum in the high-energy tail, which could be relevant for observations of UHE particles.
Figure 5: We show the fraction $N_{e}/\bar{N}_{e}$ (4.15) of accelerated electrons drifting out of bulk electrons at different temperatures $T^{+}_{\gamma}\gg m_{e}$ near the horizon. The $N_{e}/\bar{N}_{e}$ value is exponentially sensitive to the horizon temperature $T_{\gamma}^{+}$. The probability of accelerated electrons reaching high energies is small for temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, and the Compton-rocket effect and runaway dynamics might not be observationally relevant. These figures also represent the proton case by substituting the subscript $e\rightarrow p$.
For different temperatures $T^{+}_{\gamma}\approx 100m_{e}\approx 50$ MeV and $T^{+}_{\gamma}\approx 20m_{e}\approx 10$ MeV, Figure 5 shows the following spectral features of the probability that electrons get accelerated to a large Lorentz $\gamma_{e}$ factor:
  1. (a)
    for Lorentz factors $\gamma_{e}<10^{3}$, the accelerated electron fraction $N_{e}/\bar{N}_{e}$ decreases as $\gamma_{e}$ increases, which may be approximated as a decay power law
    $\frac{N_{e}}{\bar{N}_{e}}\propto(\gamma_{e})^{-\nu},~~\nu>0,~~(\gamma_{e}<10^{3});$ (4.16)
  2. (b)
    for Lorentz factors $\gamma_{e}>10^{3}$, the fraction approach a constancy
    $\frac{N_{e}}{\bar{N}_{e}}~\propto(\gamma_{e})^{0},~~~(\gamma_{e}>10^{3}).$ (4.17)
The feature (a) is expected in a usual physical situation where the probability of drifting electrons running to high energies becomes smaller, due to their kinetic energy dissipations in collisions with photons and pairs. However, the feature (b) is peculiar, showing that the probability $N_{e}/\bar{N}_{e}$ does not decrease as (a), instead, becomes very flat when $\gamma_{e}>10^{3}$. These are the exact attributes of unstable runaway behaviour. Since the KN cross-section decreases as the drifting electrons’ energies increase, they continue accelerating to higher and higher energies by the Compton-rocket effect without dissipating much energy into photons and pairs in the radiation fluid. Such an avalanche dynamics is crucially important for producing UHE particles of Lorentz factors $\gamma_{e}\gg 1$.
As the temperature $T_{\gamma}$ decreases and approaches $2m_{e}$, the asymptotic value $N_{e}/\bar{N}_{e}$ (4.17) is exponentially suppressed by many orders of magnitude. Some technical difficulties occur in making numerical calculations. Therefore, using $\Delta\gamma^{\rm KN}_{e}$ (4.12) and $(N_{e}/\bar{N}_{e})$ (4.15), we approximately calculate the asymptotic constancy (4.17) for large $\gamma_{e}>10^{3}$
@@
which is exponentially suppressed by small horizon temperatures $T^{+}_{\gamma}/m_{e}$, as can be seen in Figs. 4 and 5. This result (4.18) implies the following features. (i) The significantly nontrivial fraction $N_{e}/\bar{N}_{e}$ of high-energy electrons ($\gamma_{e}\gg 1$) can only be produced in opaque fireshells or spots of the radiation fluid of photons and pairs at temperatures $T_{\gamma}>10m_{e}$, which can be present in highly energetic arenas of astrophysical sources and the early Universe. (ii) In the case of the Sun ($T_{\odot}/m_{e}\ll 1$), the excess of large $\gamma_{e}$ electrons $N_{e}$ in the superthermal tail deviating from the blackbody radiation spectrum is too small to be relevant for observations. (iii) However, the probability of $\gamma_{e}\gg 1$ electrons’ excess may be relevant for the detections in future laser experiments [33] for the physics of strong radiation field intensities $I_{\gamma}>E^{2}_{c}/(8\pi)$, equivalently to high temperature $T^{+}_{\gamma}>10m_{e}$.
To close this section, we mention that for the optically thin system of electron-proton plasma interacting with photons at energies $\epsilon_{\gamma}\sim m_{e}$, Equations of types (4.8,4.9) or (4.13,4.14) were used in the numerical studies [27]. Their results show the superthermal spectrum tail of runaway electrons $N_{e}$ drifting backwards to massive protons due to the electric field induced by the charge separation between electrons and protons.Here, instead, we study electrons and protons that are fully ionised and immersed inside an opaque radiation fluid of photons and pairs. They interact with photons and pairs at very high temperatures $T_{\gamma}\gg m_{e}$, energy density $\rho_{\gamma}\gg m_{e}^{4}$ and number density $n_{\gamma}\gg m_{e}^{3}$.
5 Runaway protons pulled by runaway electrons
5.1 Electric field between drifting electrons and protons
A non-trivial amount of high-energy electrons running away indicates that we cannot ignore protons in the dynamics of the Compton-rocket effect on electrons, because of the electric field $E_{pe}$ developed between separated electrons and protons. This is due to the very different cross sections $\sigma_{{}_{T}}\gg\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}$ of photons scattering with electrons and protons. In the one-dimensional model, the Compton-rocket effect on protons ($p_{p}=m_{p}\beta_{p}\gamma_{p}$)
@@ 5.1 Electric field between drifting electrons and protons
is much smaller than the one (3.2) for electrons.
@@
It is a highly non-linear dynamical system. We speculate the dynamical condition when photons and pairs’ radiation force $\sigma_{{}_{T}}F_{\gamma}\propto T_{\gamma}^{4}(\beta_{e}\gamma_{e})^{2}$ in Eq. (5.3) is so strong that
  1. (i)
    a significantly non-trivial fraction $N_{e}/\bar{N}_{e}$ of accelerated electrons ($\gamma_{e}\gg 1$) drifting outward, i.e., $dp_{e}/d\tau>0$ in Eq. (5.3) for electrons;
  2. (ii)
    it thus induces a strong and local electric field $E_{pe}=eN_{e}/(4\pi r^{2})$, which is an electric field between a thin spherical capacitor of electron and proton layers in the one-dimensional model;
  3. (iii)
    this electric field binds electron and positron together, pulls massive protons $dp_{p}/d\tau\approx eE_{pe}>0$ in Eq. (5.4), where the radiation force term $\propto\sigma^{p}_{{}_{T}}$ in its right-handed side is negligible. As a result, a significantly non-trivial fraction $N_{p}/\bar{N}_{p}$ of accelerated protons ($\gamma_{p}\gg 1$) drift outwards in the radial direction, following the accelerated electrons’ fraction $N_{e}/\bar{N}_{e}$.
5.2 Strong radiation field accelerating electrons and protons
The bulk and drifting electron numbers are identical to the bulk and drifting proton numbers
@@ 5.2 Strong radiation field accelerating electrons and protons
because of the global charge neutrality of ionised electrons and protons, which are immersed inside an electrically neutral radiation fluid of photons and pairs. The combination of Eqs. (5.3) and (5.4) yields the one-dimensional acceleration equation for electrons and protons together
@@
neglecting the radiation force on protons. Under the condition of a strong enough radiation flux $F_{\gamma}$ and the electric field $E_{pe}$ between electrons and protons, we assume that drifting electrons $N_{e}$ and protons $N_{p}$ are moving together
@@
through the radiation fluid of photons and pairs. The acceleration equation (5.6) is simplified as
$(m_{e}+m_{p})\frac{d(\beta_{e}\gamma_{e})}{d\tau}$ $\approx$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1].$ (5.8)
This is the same as the electron acceleration equation (3.2) or (4.3) by introducing an effective electron mass
$m^{\rm eff}_{e}=m_{p}+m_{e}\approx m_{p},$ (5.9)
which previously appeared in Eq. (4.3) without explanations. Similarly, the combination of (3.1) and (5.1) yields the energy-gain equation for accelerated electrons and protons
@@
Equations (5.8) and (5.10) qualitatively describe an accelerating system of electrons and protons driven by the radiation force, assuming electrons and protons are tightly associated by their own electric field and accelerated to the approximately same Lorentz factor $\gamma_{e}\approx\gamma_{p}$. In this approximation, all dynamical equations remain the same as those for electrons, except for replacing the electron mass $m_{e}$ with the effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$ (5.9) for a heavy system of electrons and protons. As previously anticipated, the effective mass $m^{\rm eff}_{e}\gg m_{e}$ replaces $m_{e}$ in the dynamic equations (4.3) and (5.8). As a consequence, the large temperature $T_{\gamma}>10m_{e}$ radiation flux $F_{\gamma}$ acting on an electron is necessarily required to accelerate a heavy system of both electrons and protons.
Such a dynamical configuration should be realised in an opaque radiation fluid of photons and pairs at large temperatures $T_{\gamma}\gg m_{e}$ and number $N_{\gamma}\gg\bar{N}_{e}=\bar{N}_{p}$, where the bulk numbers $\bar{N}_{e}$ and $\bar{N}_{p}$ of fully ionised electrons and protons are small contaminations. Otherwise, if the radiation fluid temperature $T_{\gamma}$ is not large enough, instead of accelerating protons, the bulk electrons $\bar{N}_{e}$ are accelerated backwards to massive protons by the induced electric field $E_{pe}$, as numerically studied in Ref. [27].
@@
In addition, when electrons and protons drift through the radiation fluid of photons and pairs, they undergo Klein-Nishina scattering processes. The KN cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons is much smaller than the one $\sigma^{e}_{{}_{\rm KN}}$ (4.1) of drifting electrons, i.e., $\sigma^{p}_{{}_{\rm KN}}\sim{\mathcal{O}}[(m_{p}/m_{e})^{2}]\sigma^{e}_{{}_{\rm KN}}$. Therefore, we neglect the cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons in the Klein-Nishina scattering processes between an electron-proton pair and the radiation fluid of photons and pairs. In these approximations, all analyses and results in Section 4 remain the same. Both drifting electrons and protons run away, becoming UHE particles. Figures 5 also represent the fraction of accelerated protons with the substitution $e\rightarrow p$. In the following, without further specification, all discussions and calculations apply to a bound system of electrons and protons with an effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$. Note that for the case $\gamma_{e}\approx\gamma_{p}$ the accelerated protons’ energy $\gamma_{p}m_{p}$ is ${\mathcal{O}}[(m_{p}/m_{e})]$ larger than the accelerated electrons’ energy $\gamma_{e}m_{e}$.
6 Luminosities of UHE emission and blackbody radiation
@@
The accelerated electrons $N_{e}$ and protons $N_{p}$ reach ultra-high energies and stream out of the fireshell, carrying away the thermal energy $E_{\gamma}$ (2.6) of the radiation fluid of photons and pairs. We will calculate the luminosity of ultra-high-energy emissions of electrons and protons, as well as the time variation of the fireshell thermal energy $E_{\gamma}$. As will be calculated below, the luminosity and the rate $\dot{E}_{\gamma}/E_{\gamma}$ of thermal energy dissipation are small. Up to a certain timescale, the processes of thermal energy transfer and temperature cooling are nearly adiabatic, and the fireshell thermal equilibrium is maintained with slowly decreasing temperature. Therefore, we regard the fireshell as an approximate steady energy $E_{\gamma}$ reservoir, like the Sun. As a result of thermal energy dissipation and strong gravitational effect, the fireshell eventually cools down and shrinks in size.
@@ 6.1 Blackbody radiation luminosity and MeV annihilation line
Due to the large opacity and temperature gradient in the centre of the fireshell ($r_{+}<r<2r_{+}$), internal photons diffuse outwards in the radial direction. At the fireshell outer boundary $r=2r_{+}$, where the temperature is about $2m_{e}$ and pair annihilation occurs, it becomes optically thin, and photons stream out as a continuous blackbody radiation of the characteristic temperature $2m_{e}$. One obtains the blackbody luminosity by solving the radiation transfer (photon diffusion) equation [34],
@@
where the photon luminosity $L^{+}_{\gamma}=\eta L_{\rm pl}$ (2.12) and mean-free path $\xi_{\gamma}\sim 10^{-4}\lambda_{e}$ (Fig. 3) are inside the centre of the fireshell of photons and pairs. In this case, the screen factor $\xi_{\gamma}/2r_{+}\sim 8.3\times 10^{-22}$ is small for the fireshell size $\sim 2r_{+}$ (2.4). The blackbody radiation luminosity (6.1) is a constant $L^{0}_{\gamma}\sim 10^{-22}L_{\rm pl}$, and the time scale (6.2) is large. These values depend on the fireshell parameters (2.4), (2.5), (2.10), and (2.12).
This emission is analogous to the blackbody radiation from the Sun’s surface. To compare and contrast, we mention the Sun’s surface halo of the temperature $\sim 6\times 10^{3}C^{\circ}\approx 0.55$ eV and the Sun radius $6.96\times 10^{10}$cm, where electrons are not relativistic ($\langle\gamma_{e}\rangle\approx 1$) and their densities are much smaller than the fireshell under discussion. The Sun’s luminosity is $L_{\odot}=3.83\times 10^{33}{\rm ergs/sec}\approx 10^{-27}L_{\rm pl}$. Thermal photons diffuse from the Sun’s core at the temperature $T_{\odot}=15\times 10^{6}C^{0}=1.5\times 10^{4}{\rm eV}=3\times 10^{-2}m_{e}$. While in the fireshell case, the blackbody has a characteristic temperature $\sim 2m_{e}$. Moreover, the radiation spectrum contains the electron-positron annihilation line and reheating of temperature increment $(11/4)^{1/3}$ before and after annihilation around $2m_{e}\approx 1$ MeV [14].
Figure 6: For the purpose of illustrating the spectral behaviour of UHE particle luminosity, we select the parameters: the horizon temperature (2.10) $T_{\gamma}^{+}=10m_{e},20m_{e}$ and acceleration number $A=10^{-3}$ (6.4), to plot high-energy electron and proton luminosities $L_{e}/L_{\rm pl}$ (6.3) in terms of the particle Lorentz factor $\gamma_{e}$.